# Visual Agents as Fast and Slow Thinkers

### v2

@@ Abstract
Achieving human-level intelligence requires refining cognitive distinctions between System 1 and System 2 thinking. While contemporary AI, driven by large language models, demonstrates human-like traits, it falls short of genuine cognition. Transitioning from structured benchmarks to real-world scenarios presents challenges for visual agents, often leading to inaccurate and overly confident responses. To address the challenge, we introduce FaST, which incorporates the Fast and Slow Thinking mechanism into visual agents. FaST employs a switch adapter to dynamically select between System 1/2 modes, tailoring the problem-solving approach to different task complexity. It tackles uncertain and unseen objects by adjusting model confidence and integrating new contextual data. With this novel design, we advocate a flexible system, hierarchical reasoning capabilities, and a transparent decision-making pipeline, all of which contribute to its ability to emulate human-like cognitive processes in visual intelligence. Empirical results demonstrate that FaST outperforms various well-known baselines, achieving 80.8% accuracy over $VQA^{v2}$ for visual question answering and 48.7% $GIoU$ score over ReasonSeg for reasoning segmentation, demonstrate FaST’s superior performance. Extensive testing validates the efficacy and robustness of FaST’s core components, showcasing its potential to advance the development of cognitive visual agents in AI systems. The code is available at https://github.com/GuangyanS/Sys2-LLaVA.
@@ 1 Introduction
In the field of artificial intelligence, System 2 delineates a cognitive mode distinguished by deliberate, analytical, and consciously reasoned processes [1, 2, 3, 4, 5]. This mode is juxtaposed to System 1, which embodies intuitive, automatic, and unconscious cognition. Achieving human-level intelligence in AI systems necessitates the deliberate cultivation and refinement of these cognitive distinctions. This process is crucial for the development of advanced reasoning and decision-making capabilities [6, 7, 8].
@@
The emergence of foundation models marks a significant turning point, where large language model (LLM) based agents have made remarkable strides in many areas, showcasing human-like intelligence across diverse tasks [9, 10, 11]. However, this achievement is primarily attributed to some features of foundation models: overparameterization and the availability of vast, general-purpose datasets [12, 13]. It is imperative to note that while these models exhibit human-like traits ($e.g.,$ inductive and deductive reasoning [14, 15, 16]), these characteristics do not equate to the processes of System 1/2 thinking [17, 18] and are far less intelligent than human thinking.
@@
In practice, visual agents often encounter challenges when moving from controlled, structured benchmarks to complex, real-world environments [19, 11]. This difficult circumstance will result in spurious reasoning pathways, akin to hallucinations, where they struggle to acknowledge their limitations or uncertainties [20, 21]. Such an issue arises from the absence of explicit modeling of the fast and slow cognitive processes, reminiscent of human System 1 and System 2 [18, 22]. Consequently, when faced with intricate inquiries, the Multimodal Large Language Model (MLLM) frequently offers overly confident yet inaccurate responses [19, 23, 24]. Addressing this problem entails reassessing MLLM algorithms to incorporate insights from the interplay between fast and slow thinking (System 1/2) observed in human cognition. This design philosophy guides our work.
@@
In this study, we introduce the Fast and Slow Thinking (FaST) mechanism into visual agents. More concretely, we design a switch adapter to determine whether the encountered problems are best addressed using which thinking mode. Simple tasks require only fast thinking (System 1 thinking) for a straightforward problem-solving pipeline, while complex tasks necessitate the slow, deliberate processing of System 2 (see Figure. 1). Specifically, System 2 is triggered when we encounter visual challenges that have: Uncertainty: When the model has low confidence in directly identifying the object to which the complex query is referring. For example, the query asks “the appliance for storing and cooling food” instead of “refrigerator,” and Invisibility: When dealing with minuscule-sized objects that evade detection by standard visual encoders, where normal visual agents cannot tell what it is. This switch adapter is achieved by designing negative contextual data to re-adjust the model’s confidence and ignite world knowledge (see detailed analysis in §4.2). Subsequently, a proposal adapter is engaged to outline regions that are related to the questions. This allows visual agents to leverage the newly acquired data, thereby facilitating a more detailed and precise response. Further, if the inquiry necessitates detailed insights into particular instances, a seg adapter provides segmentation masks, offering additional contextual information for deeper analysis (see detailed analysis in §4.3).
@@
FaST enjoys a few attractive qualities. ❶ Flexible system: Building on a foundation that explicitly models System 1/2 thinking, our proposed method adeptly handles complex visual tasks, demonstrating competitive performance in a streamlined pipeline (see §3.2). FaST’s core epistemology combines an intuitive mechanism for straightforward cases with deliberate analytics for more intricate scenarios, thereby enhancing the development of a human-like visual agent. ❷ Hierarchical reasoning: FaST perceives visual tasks with a top-down granularity, encompassing image-level cues, box-level candidates, and pixel-level targets (see Fig. 2). This progressive approach facilitates a sensible understanding of visual content, starting from global concepts, progressing through region-specific candidate assessment, and culminating in precise target identification. Each stage of this process involves developing concrete ideas and establishing a coherent “chain of evidence” to support the final inference. ❸ Transparent pipeline: FaST’s decision-making process embodies a neuro-symbolic essence, yielding intermediate step outputs as readily interpretable symbols ($e.g.$, RoI-driven boxes or masks), facilitating direct visual inspection by humans. This inherent reasoning mechanism enables ad-hoc explainability of the model’s behavior (see Fig. 3), distinguishing FaST from prior approaches [25, 26] that lack precise explication of their operational mechanisms regularly.
@@
We conducted a series of experiments to validate the efficacy of our proposed method. In §4.1.1, we apply FaST to visual question answering and multimodal benchmarks. FaST demonstrates significantly improved performance over baselines such as LLaVA-v1.5 [26], achieving performance gains on benchmarks like TextVQA [27] with a 2.5% increase in accuracy and a total score improvement of 6.7 on MME [28]. In §4.1.2, we explore the versatility of our approach through its application to tasks such as referring and reasoning segmentation, with performance gains including an increase of 4.1% $CIoU$ with LLaVA-v1.5, and improvements of 3.2% $CIoU$ and 2.7% $GIoU$ on the ReasonSeg dataset over LISA-7B [29]. The robustness and effectiveness of the core components of our framework are further substantiated through a series of ablation studies, as elaborated in §4.3.
2 Related Work
LLM as Visual Agents. With the capabilities that LLMs have demonstrated in language understanding and generation [30, 13, 31, 32, 33], the research community has progressed to explore how LLMs can be enhanced with vision input for multimodal tasks as visual agents [34, 35, 36, 11, 37, 25, 26]. There are two paradigms for LLM-based visual agents: end-to-end based and tool-using visual agents. Following the principle of instruction tuning, end-to-end visual agents are trained with a curated visual instruction tuning dataset to digest features from multi-modality, unlocking the capability to answer visual questions [38, 39, 40, 41, 42, 43, 44, 45, 27]. For other visual tasks ($e.g.$, Segmentation, Detection, etc), end-to-end trained tailored agents can further perform downstream tasks [46, 47, 29, 48, 49, 37, 50, 51, 52, 53]. Recent research has focused on leveraging improved vision encoders and fostering more detailed visual understanding, yielding promising results [54, 55, 56]. While these approaches can be implemented with direct instruction tuning data, they represent a ‘System 1’ type of training. This type of training primarily relies on the dataset’s quality and tends to provide direct answers that are prone to hallucinations, a consequence inherent to the nature of System 1 instruction tuning data. For the second paradigm, tool-using models are built on top of a frozen LLM with access to pretrained visual perception tools [57, 58, 59]. In this scenario, the LLM first selects visual tools and then decides by thoroughly analyzing the fine-grained information extracted by visual tools [59, 60, 61]. Although the utilization of external visual tools increases the interpretability of the reasoning process, inaccuracies can arise in light of the system’s complexity. In addition, due to the complexity of reasoning, the vast amount of information generated by the agent may overwhelm key ${}_{\!}$information ${}_{\!}$related to the question query, ${}_{\!}$leading to incorrect responses.
System 2 in AI. Recently, LLMs have been engineered to produce text that mimics the step-by-step reasoning process characteristic of human cognition, akin to the analytical and deliberate thought processes associated with what is termed as System 2 in the human cognition process [62, 63, 64, 65, 66]. The systematic approach to problem-solving is a hallmark across various domains, including mathematical word problems [10, 2, 67, 68, 69, 70, 71], logical reasoning [72, 73, 74, 75, 76, 77, 16], and multi-modal reasoning [78, 60, 19]. In Explainable AI, this systematic method is emulated by the model as it generates a text-based elucidation of its reasoning and decision-making process through step by step reasoning process ($e.g.$, chain of thought) [79, 80, 81, 5]. However, it is crucial to recognize that LLMs, while powerful, are not exempt from encountering challenges when facing complex problems. One such challenge is the issue of hallucination [82, 83, 84, 85, 86, 87], which can distort the model’s reasoning process and lead to inaccuracies in the explanations provided. Initially, LLM reasoning is seen as only a linear chain of thoughts, where each step in the reasoning process is clearly articulated. As models evolve, they adopt more complex structures like hierarchical trees [88, 18] and intricate graphs [74], which enable them to handle much more complex problems but also restrict their general applicability because of increased topological complexity [18, 73, 74, 76, 75]. Moreover, these complex structures can lead to errors that propagate through the model’s reasoning, causing a cascade of mistakes [89]. To counter this, incorporating feedback from intermediate reasoning steps and employing iterative refinement, which is similar to human reflection, could help mitigate errors [90, 91, 92, 93, 94, 19]. In unsupervised scenarios, such feedback is vital for enhancing the reasoning capabilities of LLMs and reducing errors [95].
Our key contribution is the introduction of the chain of evidence within multimodal reasoning frameworks. This methodology enriches each reasoning step with accurate, image-based cascading information, effectively mirroring human visual and System 2 cognitive processes. Our approach not only enhances accuracy but also significantly improves interpretability and generalization capabilities.
3 Methodology
3.1 Preliminary
@@
Notation. The integration of components in visual agents $\mathcal{F}$ (based on the Large Language Model) typically involves a visual encoder, denoted as $\mathcal{E}_{V}$, a nature language encoder, represented by $\mathcal{E}_{L}$, and a language model such as Vicuna [96]. Initially, the visual agent is presented with an image $\mathcal{I}$ and an accompanying textual prompt $\mathcal{Q}$, which could be a question or instruction. Then the visual agent combines these multimodal tokens into a united space. Finally, the visual agent outputs a textual response $R$ given the textual and image input. The generation process can be expressed as Eq. 1:
@@
(System 1 and System 2) System 1 and System 2 are two different systems of thinking proposed by Nobel Laureate Daniel Kahneman in his book Thinking, Fast and Slow [22].
System 1 (Fast Thinking System): Unconscious, automated thinking processes, fast, intuitive, effortless, responsible for automatic responses and basic cognitive operations in daily activities, vulnerable to heuristic biases and errors, $e.g.$, recognizing familiar faces, and knowing the location of objects.
System 2 (Slow Thinking System): Conscious, energetic thinking processes, slow, effortful, logical, and analytical, responsible for complex calculations, reasoning, and decision-making, can monitor and control System 1 processes, $e.g.$ filling out a tax form, finding the position of a word in a sentence.
@@ 3.2 FaST
(Switching-friendly dataset) A Negative Data for Target Objects Reasoning Dataset $\mathcal{D}$ of 100,000 (image, question, answer) triples was constructed to facilitate the identification of target regions or objects required to answer a question. The dataset constructs questions about the absence or details of certain objects, deliberately made too small to be perceived by the visual encoder.
@@
(Switch Adapter) A light-weight adapter that is fine-tuned with both positive fast-thinking data and negative data (Def. 3.1) to acquire system switching capability. When the adapter encounters harder questions, the switch mechanism will be triggered for later slow thinking.
@@
Specifically, we use negative data that contain missing objects $\mathcal{O}_{missing}$ and context clues $\mathcal{C}_{clue}$ for training the system switch adapter for triggering the slow mode as Def. 3.2. The slow mode should deal with question-image pairs that are 1) uncertain in pinpointing the specific object in question, and 2) too small to perceive for the standard visual encoder. So we utilize triplet data in dataset Def. 3.1 (image, question, answer) as where the question requires objects that are not in the image or too small to be perceived by the visual encoder. The threshold is set to be $20\times 20$. We require the model to tell that certain objects are missing instead of a direct answer, and we utilize the world knowledge to list all the objects and also the context information for later deliberate reasoning.
@@
$[Bboxes]=\mathcal{P}_{adapter}\left[\mathcal{E}_{L}(Q),Region,\mathcal{C}_{clue}\right]$ (5)
@@
(Pixel-level mask decoder) The Pixel-level mask decoder is the decoder of segment anything(SAM [97]). The pixel-level mask decoder is fine-tuned to produce target masks based on the hidden embeddings.
When we have a more specific target proposal(bounding box $[Bboxes]$), FaST will apply a fine-grained pixel-level mask decoder $\mathcal{P}_{Seg}$ as Eq. 3.4 to output the specific mask part $[Mask]$ of the target proposal $[Bboxes]$ to focus on as Eq. 6. We name this whole process from $Region$ to $[Mask]$ chain of evidence as Def. 3.5 similar to thinking more and more deeply by humans.
@@
The decision-making process inherent in FaST is distinguished by its neuro-symbolic nature, which generates intermediate outputs as easily interpretable symbols, including region-of-interest (RoI) driven boxes and object-driven masks. This capability allows for direct visual inspection by humans, thereby augmenting the transparency of the model’s operations. Moreover, the intrinsic reasoning mechanism of FaST enhances ad-hoc explainability of its behavior, shown in Fig. 2.
@@ 3.3 Implementation Details
The framework of FaST as Fig. 2’s implementation details and in are shown in this section below.
$\bullet$ Visual Agents. We choose the architecture and configuration of LLaVA-v1.5 [26] as our visual agent. The most important component in a visual agent is the visual encoder $\mathcal{E}_{V}(I)$: A CLIP-ViT-L-336px model [98] is used, where input images are resized or padded to $336*336$ pixels, learning to associate visual features with corresponding textual descriptions. An MLP projection with channels of $[256,4096,4096]$ is used for connecting image representations into the word embedding space.
$\bullet$ Mask Decoder. The mask decoder $\mathcal{P}_{seg}$ architecture is identical to SAM. Besides, it is fully fine-tuned with a collection of semantic segmentation [99, 100, 101, 102, 103] and referring segmentation [104, 105] datasets to efficiently map the <seg> token representations to a mask if the FaST need to segment.
$\bullet$ Chain of Evidence. When we apply the chain of evidence as 3.5 in the large language model to get the answer as the final step like Eq. 7. The whole sequence of the chain of evidence is too long to load in the $\mathcal{F}_{Slow}$. So FaST needs a visual sampler based on cross-attention that is trained to decrease the number of image tokens to a suitable length (from 256 to 32), apart from MLP projection.

### v3

@@ Abstract
Achieving human-level intelligence requires refining cognitive distinctions between System 1 and System 2 thinking. While contemporary AI, driven by large language models, demonstrates human-like traits, it falls short of genuine cognition. Transitioning from structured benchmarks to real-world scenarios presents challenges for visual agents, often leading to inaccurate and overly confident responses. To address the challenge, we introduce FaST, which incorporates the Fast and Slow Thinking mechanism into visual agents. FaST employs a switch adapter to dynamically select between System 1/2 modes, tailoring the problem-solving approach to different task complexity. It tackles uncertain and unseen objects by adjusting model confidence and integrating new contextual data. With this novel design, we advocate a flexible system, hierarchical reasoning capabilities, and a transparent decision-making pipeline, all of which contribute to its ability to emulate human-like cognitive processes in visual intelligence. Empirical results demonstrate that FaST outperforms various well-known baselines, achieving 80.8% accuracy over $VQA^{v2}$ for visual question answering and 48.7% $GIoU$ score over ReasonSeg for reasoning segmentation, demonstrate FaST’s superior performance. Extensive testing validates the efficacy and robustness of FaST’s core components, showcasing its potential to advance the development of cognitive visual agents in AI systems. The code is available at this link.
@@ 1 Introduction
In the field of artificial intelligence, System 2 delineates a cognitive mode distinguished by deliberate, analytical, and consciously reasoned processes (Wei et al., 2022; Wang et al., 2023d; Zelikman et al., 2022; Zhou et al., 2023; Hua & Zhang, 2022). This mode is juxtaposed to System 1, which embodies intuitive, automatic, and unconscious cognition. Achieving human-level intelligence in AI systems necessitates the deliberate cultivation and refinement of these cognitive distinctions. This process is crucial for the development of advanced reasoning and decision-making capabilities (Zhang et al., 2023d; e; Hao et al., 2023).
@@
The emergence of foundation models marks a significant turning point, where Large Language Models (LLMs) based agents have made remarkable strides in many areas, showcasing human-like intelligence across diverse tasks (Brown et al., 2020; Kojima et al., 2022; Ge et al., 2023). However, this achievement is primarily attributed to some features of foundation models: overparameterization and the availability of vast, general-purpose datasets (Kaplan et al., 2020; OpenAI, 2024). It is imperative to note that while these models exhibit human-like traits ($e.g.,$ inductive and deductive reasoning (Huang & Chang, 2023b; Dasgupta et al., 2022; Jin et al., 2024b)), these characteristics do not equate to the processes of System 1/2 thinking (Nye et al., 2021; Yao et al., 2023b) and are far less intelligent than human thinking.
@@
In practice, visual agents often encounter challenges when moving from controlled, structured benchmarks to complex, real-world environments (Wu & Xie, 2024; Ge et al., 2023). This problematic circumstance will result in spurious reasoning pathways, akin to hallucinations, where they struggle to acknowledge their limitations or uncertainties (Gunjal et al., 2024; Chen et al., 2023c). Such an issue arises from the absence of explicit modeling of the fast and slow cognitive processes, reminiscent of human System 1 and System 2 thinking (Yao et al., 2023b; Kahneman, 2011). Consequently, when faced with intricate inquiries, the Multimodal Large Language Model (MLLM) frequently offers overly confident yet inaccurate responses (Wu & Xie, 2024; Chen et al., 2024b; Tong et al., 2024a). Addressing this problem entails reassessing MLLM algorithms to incorporate insights from the interplay between fast and slow thinking (System 1 and System2) observed in human cognition. Our design philosophy guides us to incorporate human qualities into our work.
@@
In this study, we introduce the Fast and Slow Thinking (FaST) mechanism into visual agents. More concretely, we design a switch adapter to determine whether the encountered problems are best addressed using which thinking mode. Simple tasks require only fast thinking (System 1) for a straightforward problem-solving pipeline, while complex tasks necessitate the slow, deliberate processing of System 2 (see Fig. 1). Specifically, System 2 is triggered when we encounter visual challenges that have: Uncertainty: When the model has low confidence in directly identifying the object to which the complex query is referring. For example, the query asks “the appliance for storing and cooling food” instead of “refrigerator,” and Invisibility: When dealing with minuscule-sized objects that evade detection by standard visual encoders, where normal visual agents cannot tell what it is. This switch adapter is achieved by designing negative contextual data to re-adjust the model’s confidence and ignite world knowledge (as detailed analysis in §3.2). Subsequently, a proposal adapter is engaged to outline regions that are related to the questions. This allows visual agents to leverage the newly acquired data, thereby facilitating a more detailed and precise response. Further, if the inquiry necessitates detailed insights into particular instances, a seg adapter provides segmentation masks, offering additional contextual information for deeper analysis (as detailed analysis in §3.3).
@@
FaST enjoys a few attractive qualities. ❶ Flexible system: Building on a foundation that explicitly models System 1/2 thinking, our proposed method adeptly handles complex visual tasks, demonstrating competitive performance in a streamlined pipeline (see §2.1). FaST’s core epistemology combines an intuitive mechanism for straightforward cases with deliberate analytics for more intricate scenarios, thereby enhancing the development of a human-like visual agent. ❷ Hierarchical reasoning: FaST perceives visual tasks with a top-down granularity, encompassing image-level cues, box-level candidates, and pixel-level targets (see Fig. 2). This progressive approach facilitates a sensible understanding of visual content, starting from global concepts, progressing through region-specific candidate assessment, and culminating in precise target identification. Each stage involves developing concrete ideas and establishing a coherent “chain of evidence” to support the final inference. ❸ Transparent pipeline: FaST’s decision-making process embodies a neuro-symbolic essence in System 2 mode, yielding intermediate step outputs as interpretable symbols ($e.g.$, bounding boxes or masks), facilitating direct visual inspection by humans. This inherent reasoning mechanism enables ad-hoc explainability of the model’s behavior (see Fig. 3), distinguishing FaST from prior approaches (Liu et al., 2024a) that lack precise explication of their operational mechanisms.
@@
We conducted a series of experiments to validate the efficacy of our proposed method. In §3.1.1, we apply FaST to visual question answering and multimodal benchmarks. FaST demonstrates significantly improved performance over baselines such as LLaVA-v1.5 (Liu et al., 2024a), achieving performance gains on benchmarks like TextVQA (Singh et al., 2019) with a 2.5% increase in accuracy and a total score improvement of 6.7 on MME (Fu et al., 2024). In §3.1.2, we explore the versatility of our approach through its application to tasks such as referring and reasoning segmentation, with performance gains including an increase of 4.1% $CIoU$ with LLaVA-v1.5, and improvements of 3.2% $CIoU$ and 2.7% $GIoU$ on the ReasonSeg dataset over LISA-7B (Lai et al., 2024). The robustness and effectiveness of the core components of our FaST framework are further substantiated through a series of ablation studies, as elaborated in §3.3.
2 Methods
@@
Notation. The integration of components in visual agents $\mathcal{F}$ (based on the Large Language Model) typically involves a visual encoder, denoted as $\mathcal{E}_{V}$, a nature language encoder, represented by $\mathcal{E}_{L}$, and a Language Language Model such as Vicuna (Chiang et al., 2023). Initially, the visual agent is presented with an image $\mathcal{I}$ and an accompanying textual prompt $\mathcal{Q}$, which could be a question or instruction. Then the visual agent combines these multimodal tokens into a united space. Finally, the visual agent outputs a textual response $R$ given the textual and image input. The generation process can be expressed as Eq. 1:
@@
(System 1 and System 2) System 1 and 2 are two different systems of thinking proposed by Nobel Laureate Daniel Kahneman in his book Thinking, Fast and Slow (Kahneman, 2011).
System 1 (Fast Thinking): Unconscious, automated thinking processes, fast, intuitive, effortless responsible for automatic responses and basic cognitive operations in daily activities, vulnerable to heuristic biases and errors, $e.g.$, recognizing familiar faces, and knowing the location of objects.
System 2 (Slow Thinking): Conscious, energetic thinking processes, slow, effortful, logical, and analytical, responsible for complex calculations, reasoning, and decision-making, can monitor and control System 1 processes, $e.g.$ filling out a tax form, finding the position of a word in a sentence.
@@ 2.1 FaST
(Switching-friendly dataset) A Negative Data for Target Objects Reasoning Dataset $\mathcal{D}$ of 100,000 (image, question, answer) triples was constructed to facilitate the identification of target regions or objects required to answer a question. The dataset constructs questions about the absence or details of certain objects, deliberately made too small to be perceived by the visual encoder. Section A for more details.
@@
(Switch Adapter) A light-weight adapter that is fine-tuned with both positive fast-thinking data and negative data (Remark. 2.1) to acquire system switching capability. When the adapter encounters harder questions, the switch mechanism will be triggered for later slow thinking.
@@
Specifically, we use negative data that contain missing objects $\mathcal{O}_{missing}$ and context clues $\mathcal{C}_{clue}$ for training the system switch adapter for triggering the slow mode as Remark. 2.2. The slow mode should deal with question-image pairs that are 1) uncertain in pinpointing the specific object in question, and 2) too small to perceive for the standard visual encoder. So we utilize triplet data in dataset Remark. 2.1 (image, question, answer) as where the question requires objects that are not in the image or too small to be perceived by the visual encoder. The threshold is set to be $20\times 20$. We require the model to tell that certain objects are missing instead of a direct answer, and we utilize the world knowledge to list all the objects and also the context information for later deliberate reasoning.
2.2 Preliminary
@@
[…]
@@
(Pixel-level mask decoder) The Pixel-level mask decoder is the decoder of segment anything(SAM (Kirillov et al., 2023)). The pixel-level mask decoder is fine-tuned to produce target masks based on the hidden embeddings.
When we have a more specific target proposal(bounding box $[Bboxes]$), FaST will apply a fine-grained pixel-level mask decoder $\mathcal{P}_{Seg}$ as Eq. 6 to output the specific mask part $[Mask]$ of the target proposal $[Bboxes]$ to focus on as Eq. 6. We name this whole process from $Region$ to $[Mask]$ chain of evidence as Remark. 2.5 similar to thinking more and more deeply by humans.
@@
The decision-making process in FaST is distinguished by its neuro-symbolic nature, which generates intermediate outputs as easily interpretable symbols, including region-of-interest (RoI) driven boxes and object-driven masks. This capability allows humans to perform direct visual inspections, thereby augmenting the transparency of the model’s operations. Moreover, the intrinsic reasoning mechanism of FaST enhances the ad-hoc explainability of its behavior, see Fig. 2.
@@ 2.3 Implementation Details
The framework of FaST(as Fig. 2)’s implementation details are shown in this section below.
$\bullet$ Visual Agents. We choose the architecture and configuration of LLaVA-v1.5 (Liu et al., 2024a) as our visual agent. The most important component in a visual agent is the visual encoder $\mathcal{E}_{V}(I)$: A CLIP-ViT-L-336px model (Radford et al., 2021) is used, where input images are resized or padded to $336*336$ pixels, learning to associate visual features with corresponding textual descriptions. An MLP projection with channels of $[256,4096,4096]$ is used for connecting image representations into the word embedding space.
$\bullet$ Mask Decoder. The mask decoder $\mathcal{P}_{seg}$ architecture is identical to SAM. Besides, it is fully fine-tuned with a collection of semantic segmentation (Caesar et al., 2018; Zhou et al., 2017; Ramanathan et al., 2023; He et al., 2022; Chen et al., 2014) and referring segmentation (Mao et al., 2016; Kazemzadeh et al., 2014) datasets to efficiently map the <seg>token representations to a mask if the FaST need to segment.
$\bullet$ Chain of Evidence. When we apply the chain of evidence as Remark 2.5 in the LLM to get the answer as the final step like Eq. 7. The whole sequence of the chain of evidence is too long to load in the $\mathcal{F}_{Slow}$. So FaST needs a visual sampler based on cross-attention that is trained to decrease the number of image tokens to a suitable length (from 256 to 32), apart from MLP projection.